THE VIBRATION-ROTATION EMISSION SPECTRA OF GASEOUS ZnH_{2} AND ZnD_{2}

\underline{ALIREZA SHAYESTEH}, DOMINIQUE R. T. APPADOO, and P. F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, ON, N2L 3G1, Canada; IOULI E. GORDON, Department of Physics, University of Waterloo, Waterloo, ON, N2L 3G1, Canada.

High resolution infrared emission spectra of gaseous ZnH_{2} and ZnD_{2} have been recorded with a Fourier transform spectrometer. The molecules were generated in an emission source that combines an electrical discharge with a high temperature furnace. The vibration-rotation emission spectra of ZnH_{2} and ZnD_{2} were recorded in the 1200-2200 cm^{−1} region at an instrumental resolution of 0.01 cm^{−1}. The antisymmetric stretching fundamental bands, 001-000, of ^{64}ZnH_{2} and 64ZnD_{2} were observed near 1889.4 cm^{−1} and 1371.6 cm^{−1}, respectively, and for the minor isotopes of zinc, ^{66}Zn and ^{68}Zn, the band origins were shifted by approximately 1-2 cm^{−1}. Preliminary analysis of the spectra resulted in r_{0} values of 1.535271(1) Å and 1.531833(9) Å for ^{64}ZnH_{2} and ^{64}ZnD_{2}, respectively. Several hot bands of ZnH_{2} involving ν_{1}, ν_{2} and ν_{3} have also been observed, and analysis of these bands will lead to an equilibrium structure (r_{e}) for ZnH_{2}. The results will be presented at the symposium.